150kg body is at rest calculate the constant force needed to impact a velocity of 8ms-1 to it in 5seconds

Grade:10

1 Answers

Mayank Sudan
2 years ago
Mass=150kg
u=0
v=8ms-1
t=5 sec
We know that ,
a=(v-u)/t
  =(8-0)/5
a=8/5
Now, 
F=M×a
  =150×8/5
  =240N
And that's the required solution.